OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (anonymous):

do you know what the equation for a line is?

OpenStudy (anonymous):

y-(-1)=-2(x-5) do you need that in y=mx+b form though or is that ok?

OpenStudy (anonymous):

thats perfect! thanks lovee bugg!

OpenStudy (anonymous):

yep no problem anything else just ask